# Adaptive Science Operations in Deep Space Missions Using Offline Belief State Planning

### v1

@@ I Introduction
Deep space exploration in extreme environments requires autonomous systems capable of maximizing science return under harsh conditions without human-in-the-loop operations. Missions such as NASA’s Europa Clipper [1] and ESA’s JUICE [2] face Earth-communication delays exceeding one hour [3], preventing timely human intervention. In astrobiological missions, biosignatures such as cell membranes can degrade within hours to days of sampling [4][5]. Testing delays risk the irretrievable loss of scientific data, potentially compromising mission objectives. This challenge stresses the need for autonomous systems capable of making timely, critical decisions onboard to maximize science return.
@@
Spacecraft autonomy has advanced across multiple domains, enhancing real-time data processing, navigation, trajectory planning, and fault detection [6]. For example, the Perseverance rover demonstrated autonomous navigation and target selection on Mars [7], and the Deep-space Autonomous Robotic Explorer (DARE) optimized trajectories for asteroid reconnaissance [8]. Despite these advancements, most existing autonomy architectures remain limited to foundational operational functions such as mobility and resource management rather than scientific decision-making. To enable the adoption of advanced autonomy in science operations, gaining the trust of scientific stakeholders is essential through efficient, transparent, and verifiable techniques that reflect expert preferences. Offline, precomputed decision policies offer a pathway to building such autonomous systems, as they allow mission operators to rigorously validate behavior prior to deployment and ensure alignment with mission objectives and constraints.
@@
To address this need, we build upon a partially observable Markov decision process (POMDP), a mathematical framework for sequential decision-making under uncertainty and partial observability. This framework has been successfully used in various space applications, including collision avoidance [9] and planetary navigation [10]. POMDPs use probabilistic observation models to maintain beliefs about uncertain environments and optimize long-term objectives.
@@
In this work, we use a POMDP framework to automate operation of the Enceladus Orbilander life detection instruments. We structure the observation model as a Bayesian network to efficiently capture complex correlations between biosignatures. An offline POMDP solver, SARSOP [11], is used to precompute optimal instrument-use policies in the form of a lookup table. These policies can be thoroughly validated prior to launch while still adapting in real time to instrument findings and off-nominal events. To evaluate our method, we compare its performance against the baseline Orbilander concept of operations using metrics such as life detection false positive rates, false negative rates, and performance in off-nominal scenarios. Our results demonstrate up to a 40% reduction in identification errors compared to the baseline, underscoring the benefits of offline belief state planning.
@@ II Enceladus Orbilander Mission Concept
We seek to automate operation of the Life Detection Suite (LDS) of the Enceladus Orbilander, a flagship mission concept from the 2023–2032 Planetary Science Decadal Survey [12]. Our goal is to develop a decision-making framework that autonomously sequences LDS instruments to maximize science return while satisfying mission constraints.
The Enceladus Orbilander, developed by NASA and the Johns Hopkins Applied Physics Laboratory (APL), is designed to assess the habitability of Enceladus and search for evidence of life. This Saturnian moon has a subsurface ocean that ejects ice material through surface plumes. Many researchers speculate these plumes contain bioactive material, making Enceladus a prime target for astrobiological exploration [4][13]. The Orbilander mission architecture combines an orbiter and a lander, passively collecting samples in a funnel during plume fly-bys and actively sampling the surface using a robotic arm scoop. To detect biosignatures, the spacecraft carries a sophisticated suite of six instruments: two mass spectrometers, an electrochemical sensor array, an organic analyzer, a microscope, and a nanopore sequencer. These instruments are capable of detecting a wide range of potential biosignatures—from amino acids and cells to weak indicators like pH (Table I). During orbit, the spacecraft is expected to collect between 1 and 20 samples depending on plume activity [14].
Several mission constraints complicate LDS operations. The spacecraft’s 12-hour orbital period limits the duration and frequency of Earth communication windows, reducing opportunities for real-time human decision-making. Power is also limited: the spacecraft’s radioisotope thermoelectric generators (RTGs) are projected to degrade by 13% before Saturn orbit insertion and up to 22% by landing [14].
@@ III Methodology
This section describes the general formulation of a POMDP [15] and how we model the Life Detection Suite of the Enceladus Orbilander within this framework. To efficiently structure sensor observations, we incorporate a Bayesian network into the observation model. We then use SARSOP [11], an approximate offline POMDP solver, to compute policies that map the vehicle’s current state and observations to the next instrument action. These policies are optimized to maximize biosignature detection accuracy while maintaining robustness under off-nominal conditions.
A partially observable Markov decision process is a mathematical framework for modeling sequential decision-making problems under uncertainty [16]. A POMDP $\mathcal{P}$ is defined as the tuple
@@ III-B Observations
POMDPs can grow quickly in complexity due to high-dimensional observation spaces, which can make solving them computationally intractable. To address this issue, we use a Bayesian network to compactly model observations. A Bayesian network is a probabilistic model that represents a set of random variables and their conditional dependencies using a directed, acyclic graph. This structure is suited to astrobiological settings, where multiple sensors may find correlated observations of the same underlying processes. By factoring the joint distribution into a product of conditional probability densities, Bayesian networks drastically reduce the number of parameters required to model samples. For example, a Bayesian network with structure $A\rightarrow B\rightarrow C$ induces a joint probability distribution that can be factored as $P(A,B,C)=P(A)\cdot P(B\mid A)\cdot P(C\mid B)$. For an overview of Bayesian networks, see [17] or [16].
@@
The Bayesian network in Figure 1 demonstrates how the existence of life $s_{L}$ directly informs the presence of various biosignatures in a sample. We employ a causal graph structure, where parent nodes directly influence the observations of their child nodes. This structure often results in sparser connections between nodes when modeling real-world systems [17]. Some nodes, $o_{1}$ through $o_{5}$, represent strong biosignatures, such as the presence of polyelectrolytes or cells, while others are weaker, more ambiguous indicators of life. For example, measured pH $o_{9}$ (not itself a biosignature) depends on amino acid abundance $o_{5}$ and polyelectrolytes $o_{1}$, which more directly depend on whether the sample contains life. This leads to a multi-layer Bayesian network structure, with the life state $s_{L}$ as the root node that affects all biochemical children nodes.
@@
For each metric in Table I, we define conditional probability distributions (CPDs) representing the likelihood of each measurement given the parent nodes. These CPDs are shown in Figure 2. Each distribution is designed using insights from existing astrobiology literature, though the underlying assumptions and parameter choices can easily be adjusted to reflect various expert opinions [4, 18, 19]. Accuracy may be limited due to the lack of empirical data on biosignatures from icy moons like Enceladus. We follow the strategies for addressing the challenges of biosignature CPD design from [19]. As future missions collect more observational data from deep space environments, both structure and CPD parameter learning can be performed to design Bayesian networks that better represent the statistical patterns of biosignature evidence. To enable exact Bayesian inference, we discretize the CPDs using histogram binning.
@@ III-F Solvers
To solve a POMDP, the agent maintains a belief $b$, a probability distribution over possible states, to track its uncertainty about the environment. This belief $b$ updates with each new observation $o$. Most solvers output action sequences that attempt to maximize expected cumulative reward under uncertainty. A mapping of an agent’s belief to an action is known as a policy, where $\pi(b)=a$. We use the SARSOP algorithm [11] to precompute an offline policy. Offline policies are static once deployed, but they can be thoroughly tested and validated prior to deployment to provide performance guarantees. SARSOP approximates the optimal value function by sampling only reachable beliefs, which improves efficiency over exact methods and is well-suited to problems with large or continuous state spaces.
@@ III-G Baseline
We compare our methodology to the Concept of Operations outlined in the Enceladus Orbilander’s mission planning document [14]. We provide the full ConOps in Algorithm 1. With each sampling cycle, the algorithm accumulates plume material up to a specified maximum volume $s_{V}^{max}$ needed to run all instruments. Once sufficient volume is collected, the algorithm sequentially activates each scientific instrument to analyze the sample. Each measurement updates the belief $b(s_{L})$. Then, if the belief exceeds the life detection threshold $T_{biotic}$, the sample is declared biotic; if it falls below the abiotic threshold $T_{abiotic}$, the absence of life is declared. This threshold-driven approach imposes a rigid structure on the life detection process, potentially limiting flexibility in interpreting ambiguous or borderline results.

### v2

@@ I Introduction
Deep space exploration in extreme environments requires autonomous systems capable of maximizing science return under harsh conditions without human-in-the-loop operations. Missions such as NASA’s Europa Clipper [undef] and ESA’s JUICE [undefa] face Earth-communication delays exceeding one hour [undefb], preventing timely human intervention. In astrobiological missions, biosignatures such as cell membranes can degrade within hours to days of sampling [undefc][undefd]. Testing delays risk the irretrievable loss of scientific data, potentially compromising mission objectives. This challenge stresses the need for autonomous systems capable of making timely, critical decisions onboard to maximize science return.
@@
Spacecraft autonomy has advanced across multiple domains, enhancing real-time data processing, navigation, trajectory planning, and fault detection [undefe]. For example, the Perseverance rover demonstrated autonomous navigation and target selection on Mars [undeff], and the Deep-space Autonomous Robotic Explorer (DARE) optimized trajectories for asteroid reconnaissance [undefg]. Despite these advancements, most existing autonomy architectures remain limited to foundational operational functions such as mobility and resource management rather than scientific decision-making. To enable the adoption of advanced autonomy in science operations, gaining the trust of scientific stakeholders is essential through efficient, transparent, and verifiable techniques that reflect expert preferences. Offline, precomputed decision policies offer a pathway to building such autonomous systems, as they allow mission operators to rigorously validate behavior prior to deployment and ensure alignment with mission objectives and constraints.
@@
To address this need, we build upon a partially observable Markov decision process (POMDP), a mathematical framework for sequential decision-making under uncertainty and partial observability. This framework has been successfully used in various space applications, including collision avoidance [undefh] and planetary navigation [undefi]. POMDPs use probabilistic observation models to maintain beliefs about uncertain environments and optimize long-term objectives.
@@
In this work, we use a POMDP framework to automate operation of the Enceladus Orbilander life detection instruments. We structure the observation model as a Bayesian network to efficiently capture complex correlations between biosignatures. An offline POMDP solver, SARSOP [undefj], is used to precompute optimal instrument-use policies in the form of a lookup table. These policies can be thoroughly validated prior to launch while still adapting in real time to instrument findings and off-nominal events. To evaluate our method, we compare its performance against the baseline Orbilander concept of operations using metrics such as life detection false positive rates, false negative rates, and performance in off-nominal scenarios. Our results demonstrate up to a 40% reduction in identification errors compared to the baseline, underscoring the benefits of offline belief state planning.
@@ II Enceladus Orbilander Mission Concept
We seek to automate operation of the Life Detection Suite (LDS) of the Enceladus Orbilander, a flagship mission concept from the 2023–2032 Planetary Science Decadal Survey [undefk]. Our goal is to develop a decision-making framework that autonomously sequences LDS instruments to maximize science return while satisfying mission constraints.
The Enceladus Orbilander, developed by NASA and the Johns Hopkins Applied Physics Laboratory (APL), is designed to assess the habitability of Enceladus and search for evidence of life. This Saturnian moon has a subsurface ocean that ejects ice material through surface plumes. Many researchers speculate these plumes contain bioactive material, making Enceladus a prime target for astrobiological exploration [undefc][undefl]. The Orbilander mission architecture combines an orbiter and a lander, passively collecting samples in a funnel during plume fly-bys and actively sampling the surface using a robotic arm scoop. To detect biosignatures, the spacecraft carries a sophisticated suite of six instruments: two mass spectrometers, an electrochemical sensor array, an organic analyzer, a microscope, and a nanopore sequencer. These instruments are capable of detecting a wide range of potential biosignatures—from amino acids and cells to weak indicators like pH (Table I). During orbit, the spacecraft is expected to collect between 1 and 20 samples depending on plume activity [undefm].
Several mission constraints complicate LDS operations. The spacecraft’s 12-hour orbital period limits the duration and frequency of Earth communication windows, reducing opportunities for real-time human decision-making. Power is also limited: the spacecraft’s radioisotope thermoelectric generators (RTGs) are projected to degrade by 13% before Saturn orbit insertion and up to 22% by landing [undefm].
@@ III Methodology
This section describes the general formulation of a POMDP [undefn] and how we model the Life Detection Suite of the Enceladus Orbilander within this framework. To efficiently structure sensor observations, we incorporate a Bayesian network into the observation model. We then use SARSOP [undefj], an approximate offline POMDP solver, to compute policies that map the vehicle’s current state and observations to the next instrument action. These policies are optimized to maximize biosignature detection accuracy while maintaining robustness under off-nominal conditions.
A partially observable Markov decision process is a mathematical framework for modeling sequential decision-making problems under uncertainty [undefo]. A POMDP $\mathcal{P}$ is defined as the tuple
@@ III-B Observations
POMDPs can grow quickly in complexity due to high-dimensional observation spaces, which can make solving them computationally intractable. To address this issue, we use a Bayesian network to compactly model observations. A Bayesian network is a probabilistic model that represents a set of random variables and their conditional dependencies using a directed, acyclic graph. This structure is suited to astrobiological settings, where multiple sensors may find correlated observations of the same underlying processes. By factoring the joint distribution into a product of conditional probability densities, Bayesian networks drastically reduce the number of parameters required to model samples. For example, a Bayesian network with structure $A\rightarrow B\rightarrow C$ induces a joint probability distribution that can be factored as $P(A,B,C)=P(A)\cdot P(B\mid A)\cdot P(C\mid B)$. For an overview of Bayesian networks, see [undefp] or [undefo].
@@
The Bayesian network in Figure 1 demonstrates how the existence of life $s_{L}$ directly informs the presence of various biosignatures in a sample. We employ a causal graph structure, where parent nodes directly influence the observations of their child nodes. This structure often results in sparser connections between nodes when modeling real-world systems [undefp]. Some nodes, $o_{1}$ through $o_{5}$, represent strong biosignatures, such as the presence of polyelectrolytes or cells, while others are weaker, more ambiguous indicators of life. For example, measured pH $o_{9}$ (not itself a biosignature) depends on amino acid abundance $o_{5}$ and polyelectrolytes $o_{1}$, which more directly depend on whether the sample contains life. This leads to a multi-layer Bayesian network structure, with the life state $s_{L}$ as the root node that affects all biochemical children nodes.
@@
For each metric in Table I, we define conditional probability distributions (CPDs) representing the likelihood of each measurement given the parent nodes. These CPDs are shown in Figure 2. Each distribution is designed using insights from existing astrobiology literature, though the underlying assumptions and parameter choices can easily be adjusted to reflect various expert opinions [undefc, undefq, undefr]. Accuracy may be limited due to the lack of empirical data on biosignatures from icy moons like Enceladus. We follow the strategies for addressing the challenges of biosignature CPD design from [undefr]. As future missions collect more observational data from deep space environments, both structure and CPD parameter learning can be performed to design Bayesian networks that better represent the statistical patterns of biosignature evidence. To enable exact Bayesian inference, we discretize the CPDs using histogram binning.
@@ III-F Solvers
To solve a POMDP, the agent maintains a belief $b$, a probability distribution over possible states, to track its uncertainty about the environment. This belief $b$ updates with each new observation $o$. Most solvers output action sequences that attempt to maximize expected cumulative reward under uncertainty. A mapping of an agent’s belief to an action is known as a policy, where $\pi(b)=a$. We use the SARSOP algorithm [undefj] to precompute an offline policy. Offline policies are static once deployed, but they can be thoroughly tested and validated prior to deployment to provide performance guarantees. SARSOP approximates the optimal value function by sampling only reachable beliefs, which improves efficiency over exact methods and is well-suited to problems with large or continuous state spaces.
@@ III-G Baseline
We compare our methodology to the Concept of Operations outlined in the Enceladus Orbilander’s mission planning document [undefm]. We provide the full ConOps in Algorithm 1. With each sampling cycle, the algorithm accumulates plume material up to a specified maximum volume $s_{V}^{max}$ needed to run all instruments. Once sufficient volume is collected, the algorithm sequentially activates each scientific instrument to analyze the sample. Each measurement updates the belief $b(s_{L})$. Then, if the belief exceeds the life detection threshold $T_{biotic}$, the sample is declared biotic; if it falls below the abiotic threshold $T_{abiotic}$, the absence of life is declared. This threshold-driven approach imposes a rigid structure on the life detection process, potentially limiting flexibility in interpreting ambiguous or borderline results.